OpenStudy (anonymous):

find the expression for the centripetal force for a body moving with a uniform speed on a circular trajectory.Help plz

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

it's centrifugal.

OpenStudy (anonymous):

@M_Vamshi_Kumar_Reddy plz help

OpenStudy (anonymous):

by newton's second law F=ma in circular motion acceleration is =v^2/r or \[\omega ^{2}r\] combining both F=m(v^2)/r or F=m(\[\omega ^{2}r\])

OpenStudy (anonymous):

@ArslanQureshi is it for centrifugal force?

OpenStudy (anonymous):

yes this formula is for centripetal force for centrifugal the formula is same but with a negative sign

OpenStudy (anonymous):

u mean f=-ma?

OpenStudy (anonymous):

yes